OpenStudy (anonymous):

Please Help me ASAP! What is the value of 7P3 ?

OpenStudy (anonymous):

P(n,r) = P(7,3) = 7! / (7 - 3)! = 210

OpenStudy (anonymous):

What is the value of 6P6?

OpenStudy (anonymous):

P(n,r) = P(6,6) = 6! / (6 - 6)! = 720